Table #10.1.6 contains the value of the house and the amount of rental income in a year that the house brings in.
a) Create a scatter plot and find a regression equation between house value and rental income.
b) Then use the regression equation to find the rental income for
a house worth $230,000 and for a house worth $400,000.
c) Find the correlation coefficient.

Data of House Value versus Rental

Value

Rental

Value

Rental

67500

6864

170000

9568

77000

4576

174000

10400

85000

7072

190000

8320

94000

8736

200000

12272

104000

7904

200000

10400

115000

7904

208000

10400

125000

7904

225000

12480

130000

9776

240000

12064

135000

7488

244500

11232

140000

9152

270000

12896

148000

8320

300000

12480

165000

13312

310000

12480
Answer #1

a)

Scatterplot 16000 14000 12000 10000 > 8000 6000 4000 2000 0 0 50000 100000 150000 200000 250000 300000 350000 X

ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 4117000.00 234000.00 109374458333 123242912 3065946000
mean 171541.67 9750.00 SSxx SSyy SSxy

Sample size,   n =   24      
here, x̅ = Σx / n=   171541.667          
ȳ = Σy/n =   9750.000          
SSxx =    Σ(x-x̅)² =    109374458333.3330      
SSxy=   Σ(x-x̅)(y-ȳ) =   3065946000.0      
              
estimated slope , ß1 = SSxy/SSxx =   3065946000/109374458333.333=   0.0280      
intercept,ß0 = y̅-ß1* x̄ =   9750- (0.028 )*171541.6667=   4941.4049      
              
Regression line is, Ŷ=   4941.405   + (   0.028   )*x

b)

Predicted Y at X=   230000   is          
Ŷ=   4941.40488   +   0.02803   *230000=   11388.68

-------------

Predicted Y at X=   400000   is          
Ŷ=   4941.40488   +   0.02803   *400000=   16154.06

c) correlation coefficient ,    r = SSxy/√(SSx.SSy) =   0.83508